please help me in this question
let the pdf f(x)=1/5 , x=1,2,3,4,5,zero elsewhere , is the plot of the discrete type of random variable X. compute E(X) and E(X^2) use these two results to find E[(x+2)^2)]

Answer by stanbon(48568)

let the pdf f(x)=1/5 , x=1,2,3,4,5,zero elsewhere , is the plot of the discrete type of random variable X. compute E(X) and E(X^2) use these two results to find E[(x+2)^2)]
---------------------
E(X)=(1/5)(1+2+3+4+5)=15/5=3
E(X^2)=(1/5)(1+4+9+16+25)=55/5=11
E(X^2+2X+4)=11+2(3)+4=24
Cheers,
Stan H.